OpenStudy (anonymous):

Which of these limits can L'Hopital's rule be applied to? lim x->0 (4cos^2x)/x lim x->0 e^(x^2)/x^2 lim x-> infinity sinx^2/x^2 lim x-> infinity (x^2-1)/(x+2) lim x-> -2 (x^2-1)/(x+2) Thank you so much for your time and help!

OpenStudy (anonymous):

Basically there are five limits here, all produce the indeterminate form just by substituting, but four can be solved through other means. One requires L'Hopital, but I'm just not sure which. There are so many that it's daunting and I'm unsure how to solve most of them. But I'll try to get started, and I would love some help!!

OpenStudy (anonymous):

All 5 can be applied with L'Hopital's Rule... Since we got a infinity/infinity situation for all 5.

OpenStudy (anonymous):

They are all indeterminite :).

OpenStudy (anonymous):

Just left alone, that is.

OpenStudy (anonymous):

Right, I know that the rule will work for all of them... But the question is multiple choice, and there is only one answer... so they are asking for which limit is the only solution L Hopital, which one uses it as a last resort. :)

ganeshie8 (ganeshie8):

lim x-> -2 (x^2-1)/(x+2) doesnt work here right ?

OpenStudy (anonymous):

I believe it would be the second one.

ganeshie8 (ganeshie8):

I dont see quite how it works here too : lim x->0 (4cos^2x)/x lim x->0 e^(x^2)/x^2

OpenStudy (anonymous):

Since the the numerator and denominator both got to 0/0 when you substitute.

OpenStudy (anonymous):

oh thank you for the observation ganeshie, let me try solving them with the rule.

ganeshie8 (ganeshie8):

cos(0) = 1 e^0 = 1

OpenStudy (anonymous):

Ahhh yea, I just realized that xD thanks Ganeshie8.

ganeshie8 (ganeshie8):

:)

OpenStudy (anonymous):

Its definitely the 4th then... infinity - constant/ infinity minus constant

ganeshie8 (ganeshie8):

yep only 3rd and 4th i think

OpenStudy (anonymous):

Noooo 4th one is okay.

ganeshie8 (ganeshie8):

Oh wait a second, 3rd wont work

ganeshie8 (ganeshie8):

Only 4th works

OpenStudy (anonymous):

Limit for the 5th one is 3/4 i believe.

OpenStudy (anonymous):

you are saying that only the fourth one will use L'Hopital?

OpenStudy (anonymous):

Yes. That is exactly what we are saying.

OpenStudy (anonymous):

I'm still a little confused, can you explain why the fourth one is the answer and why the others won't "work?"

ganeshie8 (ganeshie8):

\( \lim \limits_{x \to 0} \dfrac{4\cos^2x}{x} = \dfrac{4\cos^2 (0)}{0} = \dfrac{4*1}{0} = \dfrac{4}{0} \)

ganeshie8 (ganeshie8):

its not of form \(\dfrac{0}{0}\) or \(\dfrac{\infty}{\infty}\), so it doesnt satisfy the hypothesis requirement fo rapplying L'hospital's rule.

OpenStudy (anonymous):

oh, I had forgotten about the required indeterminate forms! Thanks so much for clearing this up!

ganeshie8 (ganeshie8):

np :) review this if u have time : http://tutorial.math.lamar.edu/Classes/CalcI/LHospitalsRule.aspx

OpenStudy (anonymous):

Always remember, if something is divided by zero in the calculus world, it doesn't always mean it is 0 ;P Just to be friendly! WELCOME TO THE CALCULUS UNIVERSEEEEE WOOOOOOH! lol

OpenStudy (anonymous):

Haha, thanks for your help, you guys! :D I'll check that out, ganeshie, sounds helpful. Question closing!

OpenStudy (anonymous):

Anytime :)